Home‎ > ‎Crew 147‎ > ‎

114-EngineeringReport

Engineering Report
01/14/2015
Rebeca Rodriguez

Resources:
Diesel: ¾ This gauge has remained steady since our arrival. Is this a normal situation?
Propane: 40%
Static tank: 15cm
Trash: 4 bags awaiting pick up

Water:
Filled up loft tank in the morning. We have continued water conservation. We look forward to your updates.

Water temperature:
The hot water has still not been attained. Please advise on additional procedures we can try.  A mitigation plan was proposed on
1/13/2015. The plan involves us running the hot water while looking at the heater display. We will go ahead with the plan as soon as
an additional water supply has been secured.

ATV:
We will check the oil level of ATV 3 and 5 on the EVA conducted tomorrow (1/15/2015). Do you have additional mitigation strategies
for us to try? We are looking forward to seeing them up and running.